OpenStudy (anonymous):

How many steps maximum does it take to perform row reduction on a 2x3 matrix? a 3x4 matrix? a 4x5 matrix? an NxN+1 matrix?

OpenStudy (mathmate):

How would you define a step? Is it in terms of the number of multiplications/divisions/additions?

OpenStudy (anonymous):

elementary row operations, addition, multiplication, etc.

OpenStudy (mathmate):

I assume a well-conditioned matrix, with strong diagonal elements, so no need to exchange rows. I also assume row reduction is done to the point where we have an upper triangular matrix with 1's on the diagonal. For an nx(n+1) matrix, we have to work on n rows. For row i, we have at the end, (n-i+2) elements. Each element has to do a reduction with each of the previous rows, so there are (i-1)*(n-i+2) multiplications for each row, and the same number of additions. The final step of each row is to divide each element by the diagonal element. There are (n-i+1) divisions to be done, excluding the diagonal, which is assumed to be 1. \[\sum_{i=1}^{n} \left( (n-i+1) \ divisions + (i-1)(n-i+2) \ multiplications\ and\ (i-1)(n-i+2)\ additions \right)\] If we assume all operations have equal weight, the the summation simplifies to: \[\sum_{i=1}^{n} \left( (n-i+1) + 2(i-1)(n-i+2) \right)\] This simplifies to: \[\sum_{i=1}^{n} \left( i(2n+5)-2i^{2}-(3+n) \right)\] which can be summed readily to: \[((n)(n-1)/2)(2n+5) - 2(n)(n+1)(2n+1)/6 + n(3+n))\] and simplified to: \[(2n^{3}+9n^{2}+n)/6\] The backward substitution can be worked out similarly.

OpenStudy (anonymous):

Wow! Thanks so much

OpenStudy (mathmate):

You're welcome!